HOW MANY OUNCES OF EACH 15% AND 25% HYDROCHLORIC ACID SOLUTIONS MUST BE COMBINED TO PRODUCE A 20 OZ OF THE SOLUTION THAT IS 20%HYDROCHLORIC ACID?
Answer by ankor@dixie-net.com(22740)

HOW MANY OUNCES OF EACH 15% AND 25% HYDROCHLORIC ACID SOLUTIONS MUST BE COMBINED TO PRODUCE A 20 OZ OF THE SOLUTION THAT IS 20%HYDROCHLORIC ACID?
:
Let x = amt of 25% acid required
the total is to be 20 oz, therefore:
(20-x) = amt of 15% acid required
:
.25x + .15(20-x) = .20(20)
.25x + 3 - .15x = 4
:
you should be able to solve this now